OpenStudy (anonymous):

Determine the missing term in the expression to make perfect square trinomial 1. x^2+3x+ _ 2. y^2-24y+_ 3. y^2- _ +64 4. X^2+ 4/7x+ _ --help me please

OpenStudy (anonymous):

remember \[(a+b)^{2} = a ^{2}+2ab +b ^{2}\] and \[(a-b)^{2} = a ^{2}-2ab +b ^{2}\] this are the two types of perfect squares illustrated here, now can you find b?

OpenStudy (anonymous):

is the b in #1 is 2?

OpenStudy (anonymous):

nope, remember, its a perfect square, so in #1 2ab =3; and a =1, so 2*1*b =3 b = 3/2 and hence \[b ^{2}\] = 9/4, got it?

OpenStudy (anonymous):

oww. got it. thanks

OpenStudy (anonymous):

last question, how will you find the B^2 when the middle term is fraction?